1 2. Assume you are trying to compare Project Alpha to Project Beta. Both projects have the same appropriate discount rate. Project Alpha generates infinite after tax cash flows which do not grow; the first cash flow is $80M one year from now (T=1) and will cost $600M today (T=0). Project Beta generates infinite after tax cash flows which do not grow; the first cash flow is $40M one year from now (T1) and will cost $300M today (T 0). Which statement about the crossover rate for Project Alpha and Proiect Beta is most accurate? a. If the appropriate discount rate for both projects was below the crossover rate of 13.3%, Project Alpha b. If the appropriate discount rate for both projects was below the crossover rate of 4.4%, Project Alpha c. If the appropriate discount rate for both projects was above the crossover rate of 4.4%, Project Alpha d. If the appropriate discount rate for both projects was above the crossover rate of 13.3%, Project Alpha e. If the appropriate discount rate for both projects was below the crossover rate of 20.0%, Project Alpha would have a better NPV than Project Beta would have a better NPV than Project Beta would have a better NPV than Project Beta would have a better NPV than Project Beta would have a better NPV than Project Beta 1 3. You owe your credit company $50,000 and plan to make no payments on your credit card for 1.0 year. Your credit card company currently charges you an APR of 20.0%. If you make efforts to improve your credit score by 100 points, your credit card company will adjust your APR by 4.0%. As the result of you efforts, how much less interest would you pay over 1.0 year? a. $1,386 b. $2,000 c. $2,356 d. $2,443 e. $3,41:2
Answer #1

Cross over rate is the rate at which NPV of two projects are equal. and

NPV = PV of Inflows - PV of Outflows

Here both the projects forms an infinite cash inflow. So PV of Inflows can be calculated as

t1 cash flow / Discount rate

So Cross over rate (say Ke) is calculated as follows

NPV of project alpha = NPV of project beta

80M/ke - 600M = 40M/ke - 300M

80M/ke - 40M/ke = - 300M + 600M

(80M-40M)/Ke = 300M

40M/Ke = 300M

Ke = 40M/300M

Cross over rate= 13.3%

By using trial and error method;

a. If the appropriate discount rate for both the projects was below the crossover rate of 13.3%, Project Alpha would have a better NPV than Project Beta.

Note: As per HOMEWORKLIB RULES, when multiple questions are posted, we are mandated do only the first question. So please post the questions individually.